OpenStudy (anonymous):

Explain how to solve: Of 90 coins in a piggy bank, 20% are quarters. What is the least possible amount of money in the bank?

OpenStudy (anonymous):

what is the smallest amount of money possible?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Let me try again What is the smallest type of "coinage", the rest of the coins are those

OpenStudy (anonymous):

i think they want the amount ....its confusing

OpenStudy (anonymous):

It is a guided problem solving problem which mean it gives you clues....brb

OpenStudy (anonymous):

You are thinking about it too much there are quarters, dimes and nickels and ???

OpenStudy (anonymous):

How many coins in the bank are quarters? what is the value of the coins that are quarters? What is the least possible value of the coins that are not quarters?

OpenStudy (anonymous):

90 coins, of those coins 20 / 100 are quarters, 80 / 100 are ???

OpenStudy (anonymous):

(90 * 20) / 100 = 9 * 2 /1 = 18 are quarters.... Which mean 90 - 18 are another type of coin

OpenStudy (anonymous):

.2 * 90 = 18 Quarters 90 - 18 = 72 coins remaining Assuming the rest of coins are pennies (we are trying to find least amount of money that could be in bank) we can find the smallest amount possible... (18 * 0.25) + (72 * .01) = $5.22

OpenStudy (anonymous):

OK THANKS :) KNOW I AM CERTAIN MY ANSWER WAS SOLVED CORRECTLY :)

OpenStudy (anonymous):

Another way to work it out would have been with 90 pieces of paper. I wanted to help and not just give the answer

OpenStudy (anonymous):

YOU DID :) THANKS :) its weird because i get the problems that seem hard to my friends but the ones where your like DUHH after you know how to solve them you dont get at all....